OpenStudy (anonymous):

Factor the trinomial: x2 - 7x - 8

OpenStudy (math&ing001):

x2 -7x - 8 = x2 - 8x + x - 8 = x(x-8) + x-8 = (x+1)(x-8)

OpenStudy (anonymous):

Thank you!

OpenStudy (math&ing001):

Welcome !